Rank the given algorithmic functions in their right order of growth from smallest to the largest. Use numbers as suggested in the table below to show the order. (1 being slowest growth and 4 being the fastest growth)

Algorithmic Function Rank in order of growth (1 for the slowest growth and 4 for the fastest growth)
?(?) = 300? + 6
?(? 2 ????)
?(????)
?(?) = 6? 2 + 1
Answer #1

ANSWER:

2

3

1

4

f(x) = 300n +6 = O(n)>2 0 (n2 logh)>3 o (login) => 1 (slowest) 0 (02) 4 672 +1= g(x) - V fastest